For 1/3x+2=2x+7, how do you write out the linear equation?

Answer by TimothyLamb(4379)

By algebraic simplification:
---
(1/3)x + 2 = 2x + 7
2x - (1/3)x = 2 - 7
2x - (1/3)x = -5
2(3/3)x - (1/3)x = -5
(6/3)x - (1/3)x = -5
(5/3)x = -5
5x = 3(-5)
x = -3
---
Answer:
The line is a vertical line with x-intercept a = -3 and equation x = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php